Leticia can assemble a swing set in 3 hours while Betty can in 5 hours. How long would it take for them to assemble the swing set together?
Answer by nerdybill(7384)

Leticia can assemble a swing set in 3 hours while Betty can in 5 hours. How long would it take for them to assemble the swing set together?
.
Let x = hours it takes together
then
x(1/3 + 1/5) = 1
multiply both sides by 15:
x(5 + 3) = 15
x(8) = 15
x = 15/8
x = 1.875 hours
or
x = 1 hour and 53 minutes